Log Book Entry No.23

Days, Calendar Created: 25 Oct 2013 12:00:08 Days, Calendar Last Revised: 25 Oct 2013 16:38:47 Person PaulHallam Text Dark Study


Finish taking dark frames for temperature of -10 degrees

Note: Error it temperature is +or- 0.2 degrees

Cooled down from T = 20degrees to T = -10 degrees within 5minutes - waited 30mins before taking data to ensure temperature was correct.

Exposure times 0.1s, 1s, 5s, 10s, 60s, 180s

Using Autosave feature to speed up process, Hence file format has changed to: T-10.0-001-E0001

Must go back to data taken on the 18/10/2013 and rename files for consistancy -> Done.

Proceeded to take another 120 dark frames at -7.5 degrees, as with coolers on 100% sensor temp would not go below -10.6 degrees.

Proceeded to take anothe 120 dark frames at -2.5 degrees.

CCD will then be reset to 0 degrees, and frames will be taken every 30 seconds as it is cooled from 0 to -10 degrees. The exposure time will be 2 seconds. Due to lack of time this will be postponed until next week.

Frames stored at:


Back to: LogBookAstronomy

Subject Dark Study
Status blank
Edit | Attach | Watch | Print version | History: r4 < r3 < r2 < r1 | Backlinks | Raw View | Raw edit | More topic actions

Physics WebpagesRHUL WebpagesCampus Connect • Royal Holloway, University of London, Egham, Surrey TW20 0EX; Tel/Fax +44 (0)1784 434455/437520

Topic revision: r4 - 25 Oct 2013 - PaulHallam

This site is powered by the TWiki collaboration platform Powered by PerlCopyright © 2008-2024 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding RHUL Physics Department TWiki? Send feedback